Find the length and width of a rectangular ice cream sandwich if its volume is 9 in cubed.
Height = 3/4
Width=x
Length= x+ 4

Answer by ewatrrr(24785)

 
Hi,

Find the length and width of a rectangular ice cream sandwich if its volume is 9 in cubed,
Height = 3/4, Width=x, Length= x+ 4
Question states******
(x+4)(x)(3/4) = 9 in^3
Solving for x
x^2 + 4x = 9(4/3)
x^2 + 4x -12 = 0
(x+6)(x-2) = 0
(x+6) = 0 , x = -6 throwing out negative solution
(x-2) = 0 , x = 2in, the width and the length is 6in
CHECKING our Answer***
(6)(2)(3/4) = 36/4 = 9 in^3